# Local analysis of iterative reconstruction from discrete generalized Radon transform data in the plane

### v1

@@ 8.1. Reduction of the DTB function to an integral
Now we study $G_{\epsilon}$ introduced in (5.7). As usual, suppose $x_{0}=0_{2}$. Change variables $\hat{\xi}=\epsilon\xi$ and $\check{x}=x/\epsilon$, $\check{z}=z/\epsilon$ in (5.7) and express $G_{\epsilon}$ as follows:
(8.1) $\begin{split}&G_{\epsilon}(\check{x}):=\frac{1}{(2\pi)^{2}}\int_{\mathcal{U}/\epsilon}\int_{\mathbb{R}^{2}}\big{(}\tilde{D}_{0}(\epsilon\check{x},\hat{\xi})+\kappa|\hat{\xi}|^{3}\big{)}^{-1}e^{-i\hat{\xi}\cdot(\check{x}-\check{z})}\text{d}\hat{\xi}\,F_{\epsilon}(\epsilon\check{z})\text{d}\check{z}.\end{split}$
For clarity, $G_{\epsilon}$ in (LABEL:to_scl_PDO_2) is different from the functions $G$ and $G_{0}$ used in sections 6 and 7. Since
@@
Our aim is to reconstruct the jump of $f$ at $x_{0}=0_{2}$, so we compute $\Delta G_{\epsilon}(\check{x}):=G_{\epsilon}(\check{x})-G_{\epsilon}(0_{2})$. By (LABEL:to_scl_PDO_2), $\Delta G_{\epsilon}(\check{x})=\Delta G_{\epsilon}^{(1)}(\check{x})+\Delta G_{\epsilon}^{(2)}(\check{x})$, where
(8.3) $\begin{split}\Delta G_{\epsilon}^{(1)}(\check{x})=\frac{1}{(2\pi)^{2}}\int_{\mathcal{U}/\epsilon}\int_{\mathbb{R}^{2}}&\big{[}(\tilde{D}_{0}(\epsilon\check{x},\hat{\xi})+\kappa|\hat{\xi}|^{3})^{-1}-(\tilde{D}_{0}(0_{2},\hat{\xi})+\kappa|\hat{\xi}|^{3})^{-1}\big{]}\\ &\times e^{-i\hat{\xi}\cdot(\check{x}-\check{z})}\text{d}\hat{\xi}\,F_{\epsilon}(\epsilon\check{z})\text{d}\check{z},\end{split}$
(8.4) $\begin{split}&\Delta G_{\epsilon}^{(2)}(\check{x})=\frac{1}{(2\pi)^{2}}\int_{\mathcal{U}/\epsilon}\int_{\mathbb{R}^{2}}\frac{e^{-i\hat{\xi}\cdot\check{x}}-1}{\tilde{D}_{0}(0_{2},\hat{\xi})+\kappa|\hat{\xi}|^{3}}e^{i\hat{\xi}\cdot\check{z}}\text{d}\hat{\xi}\,F_{\epsilon}(\epsilon\check{z})\text{d}\check{z}.\end{split}$
@@
Let $I=(a,b)$ be an interval where $-\pi<a<0<b<\pi$. Let $D(\lambda,\theta,\epsilon)\in C^{\infty}\big{(}(0,\infty)\times I\times(0,\epsilon_{0})\big{)}$ be a function which satisfies
(8.5) $\big{|}\partial_{\theta}^{l}\partial_{\lambda}^{k}D(\lambda,\theta,\epsilon)\big{|}\leq c_{k,l}(1+\lambda)^{-3},\ (\lambda,\theta,\epsilon)\in(0,\infty)\times I\times(0,\epsilon_{0}),$
for any $k=0,1,2$, $l\in\mathbb{N}_{0}$, some $c_{k,l}$, which can be independent of $\epsilon$, and some $\epsilon_{0}>0$. Define
@@
(8.7) $J(r)=\begin{cases}O(r^{-2})&\\ O(r^{-3}),&\text{if }D(\lambda=0^{+},\theta,\epsilon)\equiv 0,\end{cases}\ r\to\infty,$
uniformly in $\epsilon\in(0,\epsilon_{0})$. Moreover, if $c_{k,l}=O(\epsilon)$, then
(8.8) $J(r)=\begin{cases}O(\epsilon)O(r^{-2})\\ O(\epsilon)O(r^{-3}),&\text{if }D(\lambda=0^{+},\theta,\epsilon)\equiv 0,\end{cases}\ \epsilon\to 0,r\to\infty.$
The expression in brackets in (8.3) satisfies (8.5) with $c_{k,l}=O(\epsilon)$ uniformly in $\check{x}$ within bounded sets (see (C.7)). Here $\hat{\xi}=\lambda(\cos\theta,\sin\theta)$, $\lambda=|\hat{\xi}|$. By the top case in (8.8), the integral with respect to $\hat{\xi}$ in (8.3) is $O(\epsilon)O(|\check{z}|^{-2})$, $|\check{z}|\to\infty$. Recall that $\check{x}$ is confined to a bounded set. Additionally, the support of $F_{\epsilon}(\epsilon\check{z})$ is of size $O(1/\epsilon)$. Hence
(8.9) $|\Delta G_{\epsilon}^{(1)}(\check{x})|\leq c\epsilon\int_{|\check{z}|\leq 1/\epsilon}(1+|\check{z}|)^{-2}\text{d}\check{z}=O(\epsilon\ln(1/\epsilon)).$
The fraction in (LABEL:del_bfe_2) satisfies (8.5) uniformly in $\check{x}$ within bounded sets. Moreover, its value at $\lambda=0^{+}$ equals zero. By the bottom case in (8.7), the integral with respect to $\hat{\xi}$ is $O(|\check{z}|^{-3})$, $|\check{z}|\to\infty$. This means that the integral with respect to $\check{z}$ in (LABEL:del_bfe_2) is absolutely convergent. In both integrals we have used Lemma 6.1.
@@
(8.10) $\Delta F_{0}(\check{x})=\lim_{\epsilon\to 0}\Delta F_{\epsilon}(\epsilon\check{x}),\ |\check{x}|\leq A,$
for any $A>0$ (see (7.1) and (7.18)). Pick any $A\gg A_{0}$ and use (8.9) to obtain from (LABEL:to_scl_PDO_2), (8.3), (LABEL:del_bfe_2):
(8.11) $\begin{split}\Delta G_{\epsilon}(\check{x})=&\frac{1}{(2\pi)^{2}}\int_{|\check{z}|\leq A}\bigg{[}\int_{\mathbb{R}^{2}}\frac{e^{-i\hat{\xi}\cdot\check{x}}-1}{\tilde{D}_{0}(0_{2},\hat{\xi})+\kappa|\hat{\xi}|^{3}}e^{i\hat{\xi}\cdot\check{z}}\text{d}\hat{\xi}\bigg{]}\Delta F_{\epsilon}(\epsilon\check{z})\text{d}\check{z}\\ &+\frac{F_{\epsilon}(0)}{2\pi}\int_{\mathbb{R}^{2}}\frac{e^{-i\hat{\xi}\cdot\check{x}}-1}{\tilde{D}_{0}(0_{2},\hat{\xi})+\kappa|\hat{\xi}|^{3}}\frac{AJ_{1}(A|\hat{\xi}|)}{|\hat{\xi}|}\text{d}\hat{\xi}\\ &+O(1/A)+O(\epsilon\ln(1/\epsilon)),\end{split}$
where $J_{1}$ is the Bessel function of the first kind. By Lemma 8.1, the integral in brackets above is $O(|\check{z}|^{-3})$, $|\check{z}|\to\infty$. Use (8.10), and the Lebesgue dominated convergence theorem to take the limit as $\epsilon\to 0$ in the first integral in (8.11) to obtain:
(8.12) $\begin{split}\frac{1}{(2\pi)^{2}}\int_{|\check{z}|\leq A}\int_{\mathbb{R}^{2}}\frac{e^{-i\hat{\xi}\cdot\check{x}}-1}{\tilde{D}_{0}(0_{2},\hat{\xi})+\kappa|\hat{\xi}|^{3}}e^{i\hat{\xi}\cdot\check{z}}\text{d}\hat{\xi}\,\Delta F_{0}(\check{z})\text{d}\check{z}.\end{split}$
Since $A\gg A_{0}$ can be arbitrarily large and $AJ_{1}(A|\hat{\xi}|)/|\hat{\xi}|\to\delta(\hat{\xi})$ as $A\to\infty$, (8.11), (8.12), and Lemma 6.1 imply
(8.13) $\begin{split}\lim_{\epsilon\to 0}\Delta G_{\epsilon}(\check{x})=&\frac{1}{(2\pi)^{2}}\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}\frac{e^{-i\hat{\xi}\cdot\check{x}}-1}{\tilde{D}_{0}(0_{2},\hat{\xi})+\kappa|\hat{\xi}|^{3}}e^{i\hat{\xi}\cdot\check{z}}\text{d}\hat{\xi}\,\Delta F_{0}(\check{z})\text{d}\check{z}.\end{split}$
@@
The last integral simplifies using that $\Delta F_{0}(\check{x})=\varphi(\Phi_{x}^{\prime}\check{x})$ for some $\varphi$ (see (7.18)). Simple transformations give
(8.14) $\begin{split}\Delta G_{0}(\check{x}):=&\lim_{\epsilon\to 0}\Delta G_{\epsilon}(\check{x})\\ =&\frac{1}{2\pi}\int_{\mathbb{R}}\bigg{[}\int_{\mathbb{R}}\frac{e^{-i\lambda\vec{\Theta}\cdot\check{x}}-1}{\tilde{D}_{0}(0_{2},\vec{\Theta}_{0})+\kappa|\lambda|^{3}}e^{i\lambda t}\text{d}\lambda\bigg{]}\varphi(|\Phi_{x}^{\prime}|t)\text{d}t,\\ \varphi(t):=&C\int_{0}^{t}h_{0}(-s)\text{d}s,\ \vec{\Theta}_{0}:=\Phi_{x}^{\prime}/|\Phi_{x}^{\prime}|,\end{split}$
@@
Since $\tilde{R}(\lambda)$ is $C^{2}$ near $\lambda=0$, $R(t)$ is absolutely integrable. Combine (7.18), (8.14), and (8.15) to obtain
(8.16) $\begin{split}&\Delta G_{0}(\check{x})=C\int_{\mathbb{R}}(R(r-t)-R(-t))\int_{-\infty}^{|\Phi_{x}^{\prime}|t}h_{0}(-s)\text{d}s\text{d}t,\ r:=\vec{\Theta}_{0}\cdot\check{x}.\end{split}$
We have extended the integral with respect to $s$ to $(-\infty,|\Phi_{x}^{\prime}|t]$ because $\int(R(r-t)-R(-t))\text{d}t\equiv 0$ for all $r$. Simple transformations give
(8.17) $\begin{split}&\Delta G_{0}(\check{x})=(C/|\Phi_{x}^{\prime}|)\int_{\mathbb{R}}h_{0}(s)\int_{s}^{s+\Phi_{x}^{\prime}\check{x}}R(t/|\Phi_{x}^{\prime}|)\text{d}t\text{d}s.\end{split}$
@@
(8.18) $\tilde{D}(0_{2},\Phi_{x}^{\prime})=2\pi\frac{W^{2}|\Phi_{x}^{\prime}|}{|\Delta_{\Phi}|}\quad$ (by (C.5), (C.6))
(8.19) $K(x_{0},\alpha_{0})=|\Phi_{x}^{\prime}|W\quad$ (by (6.4))
(8.20) $Q_{\alpha\alpha}^{\prime\prime}(x_{0},\alpha_{0})=\frac{|\Delta_{\Phi}||x_{\alpha}^{\prime}|}{|\Phi_{x}^{\prime}|}\quad$ (by (D.21), (D.22))
(8.21) $\psi(y_{0})=\frac{|H_{x}^{\prime}|}{|\Phi_{x}^{\prime}|}\quad$ (by (4.7))
@@
Let $\bar{h}_{0}$ be defined the same way as $h_{0}$ in (7.16), but without the factor $K(x_{0},\alpha_{0})$. Then, by the above formulas and (8.15):
(8.23) $\begin{split}&\Delta G_{0}(\check{x})=\tilde{f}_{0}\Upsilon\big{(}\Phi_{x}^{\prime}\check{x}\big{)},\\ &\Upsilon(r)=2\pi\frac{W^{2}}{|\Delta_{\Phi}|}\int_{\mathbb{R}}\bar{h}_{0}(u)\int_{u}^{u+r}\mathcal{F}^{-1}\bigg{(}\frac{1}{2\pi\frac{W^{2}|\Phi_{x}^{\prime}|}{|\Delta_{\Phi}|}+\kappa|\lambda|^{3}}\bigg{)}(t/|\Phi_{x}^{\prime}|)\text{d}t\text{d}u.\end{split}$
@@
(8.24) $\begin{split}\Upsilon(r)=&\frac{1}{|\Phi_{x}^{\prime}|}\int_{\mathbb{R}}\bar{h}_{0}(u)\int_{u}^{u+r}\mathcal{F}^{-1}\bigg{(}\frac{1}{1+\kappa C_{1}|\lambda|^{3}}\bigg{)}(t/|\Phi_{x}^{\prime}|)\text{d}t\text{d}u,\\ C_{1}=&\frac{|\Delta_{\Phi}|}{2\pi W^{2}|\Phi_{x}^{\prime}|}.\end{split}$
@@
If there are multiple critical points $y_{l}$ from which the singularity at $(x_{0},\xi_{0})$ is visible, the analog of (5.7) becomes
(9.1) $\begin{split}&G_{\epsilon}=\text{Op}\big{(}\big{[}\textstyle\sum_{l}\tilde{D}_{l}(x,\xi/|\xi|)+\kappa|\epsilon\xi|^{3}\big{]}^{-1}\big{)}F_{\epsilon},\ x\in\mathcal{U}_{0},\\ &\tilde{D}_{l}(x_{0},\vec{\Theta}_{0}):=2\pi\frac{W^{2}(x_{0},y_{l})|\Phi_{x}^{\prime}(x_{0},\alpha_{l})|}{|\Delta_{\Phi}(x_{0},\alpha_{l})|}.\end{split}$
@@
(9.2) $\vec{\Theta}_{0}=\xi_{0}/|\xi_{0}|=H_{x}^{\prime}(x_{0})/|H_{x}^{\prime}(x_{0})|=\Phi_{x}^{\prime}(x_{0},\alpha_{l})/|\Phi_{x}^{\prime}(x_{0},\alpha_{l})|$
for all $l$, cf. the convention about the directions of $H_{x}^{\prime}$ and $\Phi_{x}^{\prime}$ stated above assumption 3.3. Therefore, we get from (8.15) that the analog of the function $R$ is
@@
The analog (LABEL:rhs_lim_5) becomes
(9.4) $\begin{split}&\Delta G_{0}(\check{x})=\tilde{f}_{0}\textstyle\sum_{l}\Upsilon_{l}\big{(}\Phi_{x}^{\prime}(x_{0},\alpha_{l})\check{x}\big{)},\\ &\Upsilon_{l}(r)=2\pi\frac{W^{2}(x_{0},y_{l})}{|\Delta_{\Phi}(x_{0},\alpha_{l})|}\int_{\mathbb{R}}\bar{h}_{l}(u)\\ &\hskip 28.45274pt\times\int_{u}^{u+r}\mathcal{F}^{-1}\left(\big{[}\textstyle\sum_{l}\tilde{D}_{l}(x_{0},\vec{\Theta}_{0})+\kappa|\lambda|^{3}\big{]}^{-1}\right)\bigg{(}\frac{t}{|\Phi_{x}^{\prime}(x_{0},\alpha_{l})|}\bigg{)}\text{d}t\text{d}u,\\ &\bar{h}_{l}(u):=\int_{\mathbb{R}}\varphi_{\alpha}(s)\varphi_{p}(\mu s\Phi_{\alpha}^{\prime}(x_{0},\alpha_{l})+u)\text{d}s.\end{split}$

### v2

@@ 8.1. Reduction of the DTB function to an integral
Now we study $\mathcal{G}_{\epsilon}$ introduced in (5.7). As usual, suppose $x_{0}=0_{2}$. Change variables $\hat{\xi}=\epsilon\xi$ and $\check{x}=x/\epsilon$, $\check{z}=z/\epsilon$ in (5.7) and express $\mathcal{G}_{\epsilon}$ as follows:
(8.1) $\begin{split}&\mathcal{G}_{\epsilon}(\check{x}):=\frac{1}{(2\pi)^{2}}\int_{\mathcal{U}/\epsilon}\int_{\mathbb{R}^{2}}\big{(}\tilde{D}_{0}(\epsilon\check{x},\hat{\xi})+\kappa|\hat{\xi}|^{3}\big{)}^{-1}e^{-i\hat{\xi}\cdot(\check{x}-\check{z})}\text{d}\hat{\xi}\,F_{\epsilon}(\epsilon\check{z})\text{d}\check{z}.\end{split}$
Since
@@
Our aim is to reconstruct the jump of $f$ at $x_{0}=0_{2}$, so we compute $\Delta\mathcal{G}_{\epsilon}(\check{x}):=\mathcal{G}_{\epsilon}(\check{x})-\mathcal{G}_{\epsilon}(0_{2})$. By (LABEL:to_scl_PDO_2), $\Delta\mathcal{G}_{\epsilon}(\check{x})=\Delta\mathcal{G}_{\epsilon}^{(1)}(\check{x})+\Delta\mathcal{G}_{\epsilon}^{(2)}(\check{x})$, where
(8.3) $\begin{split}\Delta\mathcal{G}_{\epsilon}^{(1)}(\check{x})=\frac{1}{(2\pi)^{2}}\int_{\mathcal{U}/\epsilon}\int_{\mathbb{R}^{2}}&\big{[}(\tilde{D}_{0}(\epsilon\check{x},\hat{\xi})+\kappa|\hat{\xi}|^{3})^{-1}-(\tilde{D}_{0}(0_{2},\hat{\xi})+\kappa|\hat{\xi}|^{3})^{-1}\big{]}\\ &\times e^{-i\hat{\xi}\cdot(\check{x}-\check{z})}\text{d}\hat{\xi}\,F_{\epsilon}(\epsilon\check{z})\text{d}\check{z},\end{split}$
and
(8.4) $\begin{split}&\Delta\mathcal{G}_{\epsilon}^{(2)}(\check{x})=\frac{1}{(2\pi)^{2}}\int_{\mathcal{U}/\epsilon}\int_{\mathbb{R}^{2}}\frac{e^{-i\hat{\xi}\cdot\check{x}}-1}{\tilde{D}_{0}(0_{2},\hat{\xi})+\kappa|\hat{\xi}|^{3}}e^{i\hat{\xi}\cdot\check{z}}\text{d}\hat{\xi}\,F_{\epsilon}(\epsilon\check{z})\text{d}\check{z}.\end{split}$
@@
Let $I\subset\mathcal{I}_{\alpha}$ be an interval. Let $D(\lambda,\theta,\epsilon)\in C^{\infty}\big{(}(0,\infty)\times I\times(0,\epsilon_{0})\big{)}$ be a function which satisfies
(8.5) $\big{|}\partial_{\theta}^{l}\partial_{\lambda}^{k}D(\lambda,\theta,\epsilon)\big{|}\leq c_{k,l}\epsilon^{n}(1+\lambda)^{-3},\ (\lambda,\theta,\epsilon)\in(0,\infty)\times I\times(0,\epsilon_{0}),$
for any $k=0,1,2$, $l\in\mathbb{N}_{0}$, and some $c_{k,l}$ independent of $\epsilon$, $\epsilon_{0}>0$, and $n\in\mathbb{N}_{0}$. Define
@@
(8.7) $J(r)=\epsilon^{n}\begin{cases}O(r^{-2})&\\ O(r^{-3}),&\text{if }D(\lambda=0^{+},\theta,\epsilon)\equiv 0,\end{cases}\ r\to\infty,$
uniformly in $\epsilon\in(0,\epsilon_{0})$.
The expression in brackets in (8.3) satisfies (8.5) with $n=1$ uniformly in $\check{x}$ within bounded sets (see (C.9)). Here $\hat{\xi}=\lambda(\cos\theta,\sin\theta)$, $\lambda=|\hat{\xi}|$. By the top case in (8.7) with $n=1$, the integral with respect to $\hat{\xi}$ in (8.3) is $\epsilon O(|\check{z}|^{-2})$, $|\check{z}|\to\infty$. Recall that $\check{x}$ is confined to a bounded set. Additionally, the support of $F_{\epsilon}(\epsilon\check{z})$ is of size $O(1/\epsilon)$. Hence
(8.8) $|\Delta\mathcal{G}_{\epsilon}^{(1)}(\check{x})|\leq c\epsilon\int_{|\check{z}|\leq 1/\epsilon}(1+|\check{z}|)^{-2}\text{d}\check{z}=O(\epsilon\ln(1/\epsilon)).$
The fraction in (LABEL:del_bfe_2) satisfies (8.5) with $n=0$ uniformly in $\check{x}$ within bounded sets. Moreover, its value at $\lambda=0^{+}$ equals zero. By the bottom case in (8.7) with $n=0$, the integral with respect to $\hat{\xi}$ is $O(|\check{z}|^{-3})$, $|\check{z}|\to\infty$. This means that the integral with respect to $\check{z}$ in (LABEL:del_bfe_2) is absolutely convergent and uniformly bounded for $0<\epsilon\ll 1$. In both integrals we have used Lemma 6.1.
@@
(8.9) $\Delta F_{0}(\check{x})=\lim_{\epsilon\to 0}\Delta F_{\epsilon}(\epsilon\check{x}),\ |\check{x}|\leq A_{1},$
for any $A_{1}>0$ (see (7.3) and (7.20)). Pick any $A_{1}\gg 1$ and use (8.8) to obtain from (LABEL:to_scl_PDO_2), (8.3), (LABEL:del_bfe_2):
(8.10) $\begin{split}\Delta\mathcal{G}_{\epsilon}(\check{x})=&\frac{1}{(2\pi)^{2}}\int_{|\check{z}|\leq A_{1}}\bigg{[}\int_{\mathbb{R}^{2}}\frac{e^{-i\hat{\xi}\cdot\check{x}}-1}{\tilde{D}_{0}(0_{2},\hat{\xi})+\kappa|\hat{\xi}|^{3}}e^{i\hat{\xi}\cdot\check{z}}\text{d}\hat{\xi}\bigg{]}\Delta F_{\epsilon}(\epsilon\check{z})\text{d}\check{z}\\ &+\frac{F_{\epsilon}(0)}{(2\pi)^{2}}\int_{\mathbb{R}^{2}}\frac{e^{-i\hat{\xi}\cdot\check{x}}-1}{\tilde{D}_{0}(0_{2},\hat{\xi})+\kappa|\hat{\xi}|^{3}}\tilde{\vartheta}_{A_{1}}(\hat{\xi})\text{d}\hat{\xi}\\ &+O(1/A_{1})+O(\epsilon\ln(1/\epsilon)),\end{split}$
where $\tilde{\vartheta}_{A_{1}}(\hat{\xi})$ is the Fourier transform of the characteristic function of the ball $\{|\check{x}|\leq A_{1}\}$. By Lemma 8.1 with $n=0$, the integral in brackets above is $O(|\check{z}|^{-3})$, $|\check{z}|\to\infty$. Use (8.9), and the Lebesgue dominated convergence theorem to take the limit as $\epsilon\to 0$ in the first integral in (8.10) to obtain:
(8.11) $\begin{split}\frac{1}{(2\pi)^{2}}\int_{|\check{z}|\leq A_{1}}\int_{\mathbb{R}^{2}}\frac{e^{-i\hat{\xi}\cdot\check{x}}-1}{\tilde{D}_{0}(0_{2},\hat{\xi})+\kappa|\hat{\xi}|^{3}}e^{i\hat{\xi}\cdot\check{z}}\text{d}\hat{\xi}\,\Delta F_{0}(\check{z})\text{d}\check{z}.\end{split}$
Since $A_{1}\gg 1$ can be arbitrarily large and $\tilde{\vartheta}_{A_{1}}(\hat{\xi})\to\delta(\hat{\xi})$ as $A_{1}\to\infty$, (8.10), (8.11), and Lemma 6.1 imply
(8.12) $\begin{split}\lim_{\epsilon\to 0}\Delta\mathcal{G}_{\epsilon}(\check{x})=&\frac{1}{(2\pi)^{2}}\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}\frac{e^{-i\hat{\xi}\cdot\check{x}}-1}{\tilde{D}_{0}(0_{2},\hat{\xi})+\kappa|\hat{\xi}|^{3}}e^{i\hat{\xi}\cdot\check{z}}\text{d}\hat{\xi}\,\Delta F_{0}(\check{z})\text{d}\check{z}.\end{split}$
@@
The last integral simplifies using that $\Delta F_{0}(\check{x})=\varphi(\text{d}_{x}\Phi\cdot\check{x})$ for some $\varphi$ (see (7.20)). Simple transformations give
(8.13) $\begin{split}\Delta\mathcal{G}_{0}(\check{x}):=&\lim_{\epsilon\to 0}\Delta\mathcal{G}_{\epsilon}(\check{x})\\ =&\frac{1}{2\pi}\int_{\mathbb{R}}\bigg{[}\int_{\mathbb{R}}\frac{e^{-i\lambda\vec{\Theta}\cdot\check{x}}-1}{\tilde{D}_{0}(0_{2},\vec{\Theta}_{0})+\kappa|\lambda|^{3}}e^{i\lambda t}\text{d}\lambda\bigg{]}\varphi(|\text{d}_{x}\Phi|t)\text{d}t,\\ \varphi(t):=&C\int_{0}^{t}h_{0}(-s)\text{d}s,\ \vec{\Theta}_{0}:=\text{d}_{x}\Phi/|\text{d}_{x}\Phi|,\end{split}$
@@
Since $\tilde{R}(\lambda)$ is $C^{2}$ near $\lambda=0$ (and smooth otherwise), $R(t)$ is absolutely integrable. Combine (7.20), (8.13), and (8.14) to obtain
(8.15) $\begin{split}&\Delta\mathcal{G}_{0}(\check{x})=C\int_{\mathbb{R}}(R(r-t)-R(-t))\int_{-\infty}^{|\text{d}_{x}\Phi|t}h_{0}(-s)\text{d}s\text{d}t,\ r:=\vec{\Theta}_{0}\cdot\check{x}.\end{split}$
We have extended the integral with respect to $s$ to $(-\infty,|\text{d}_{x}\Phi|t)$ because $\int(R(r-t)-R(-t))\text{d}t\equiv 0$ for all $r$. Transforming (LABEL:rhs_lim_4) further gives
(8.16) $\begin{split}&\Delta\mathcal{G}_{0}(\check{x})=(C/|\text{d}_{x}\Phi|)\int_{\mathbb{R}}h_{0}(s)\int_{s}^{s+\text{d}_{x}\Phi\cdot\check{x}}R(t/|\text{d}_{x}\Phi|)\text{d}t\text{d}s.\end{split}$
@@
(8.17) $\tilde{D}(0_{2},\text{d}_{x}\Phi)=2\pi\frac{W^{2}|\text{d}_{x}\Phi|}{|\Delta_{\Phi}|}\quad$ (by (C.7), (C.8))
(8.18) $K(x_{0},\alpha_{0})=|\text{d}_{x}\Phi|W\quad$ (by (6.3))
(8.19) $Q_{\alpha\alpha}^{\prime\prime}(x_{0},\alpha_{0})=\frac{|\Delta_{\Phi}||x_{\alpha}^{\prime}|}{|\text{d}_{x}\Phi|}\quad$ (by (D.23), (D.24))
(8.20) $\psi(y_{0})=\frac{|\text{d}_{x}H|}{|\text{d}_{x}\Phi|}\quad$ (by (4.7))
@@
Let $\bar{h}_{0}$ be defined the same way as $h_{0}$ in (7.18), but without the factor $K(x_{0},\alpha_{0})$. By (8.14) and (LABEL:rhs_lim_4_v2)–(8.21):
(8.22) $\begin{split}&\Delta\mathcal{G}_{0}(\check{x})=\Delta f\Upsilon\big{(}\text{d}_{x}\Phi\cdot\check{x}\big{)},\\ &\Upsilon(r)=2\pi\frac{W^{2}}{|\Delta_{\Phi}|}\int_{\mathbb{R}}\bar{h}_{0}(u)\int_{u}^{u+r}\mathcal{F}^{-1}\bigg{(}\frac{1}{2\pi\frac{W^{2}|\text{d}_{x}\Phi|}{|\Delta_{\Phi}|}+\kappa|\lambda|^{3}}\bigg{)}(t/|\text{d}_{x}\Phi|)\text{d}t\text{d}u.\end{split}$
@@
(8.23) $\begin{split}\Upsilon(r)=&\frac{1}{|\text{d}_{x}\Phi|}\int_{\mathbb{R}}\bar{h}_{0}(u)\int_{u}^{u+r}\mathcal{F}^{-1}\bigg{(}\frac{1}{1+\kappa C_{1}|\lambda|^{3}}\bigg{)}(t/|\text{d}_{x}\Phi|)\text{d}t\text{d}u,\\ C_{1}=&\frac{|\Delta_{\Phi}|}{2\pi W^{2}|\text{d}_{x}\Phi|}.\end{split}$
@@
Truncated-view artifacts are reasonably low. This will likely allow for successful image segmentation and motion estimation.
@@
Suppose the GRT data are redundant and there are multiple points $y_{l}$ from which the singularity at $(x_{0},\xi_{0})$ is visible. This means that for all $(x,\xi)$ in a conic neighborhood of $(x_{0},\xi_{0})$ there are multiple local solutions $\nu_{l}(x,\xi)$ and $\alpha_{l}(x,\xi)$ to the equation $\xi=-\nu\text{d}_{x}\Phi(x,\alpha)$. The local uniqueness and smoothness of the solutions follows from Assumption 3.2(1). By Assumption 3.3(2), their number is uniformly bounded. Set $p_{l}(x,\xi)=\Phi(x,\alpha_{l}(x,\xi))$ and $y_{l}=(\alpha_{l},p_{l})$. The principal symbol of $\mathcal{R}^{*}\mathcal{R}$ is given by (cf. (C.6))
(9.1) $\begin{split}\textstyle\sum_{l}\frac{2\pi}{|\nu_{l}|}\frac{W^{2}(x,y_{l})}{|\Delta_{\Phi}(x,\alpha)|},\ y_{l}=y(x,\xi),\nu_{l}=\nu(x,\xi).\end{split}$
Similarly to (C.8), define
(9.2) $\begin{split}\tilde{D}_{0}(x,\xi):=&2\pi\textstyle\sum_{l}\frac{|\xi|}{|\nu_{l}|}\frac{W^{2}(x,y_{l})}{|\Delta_{\Phi}(x,\alpha_{l})|},\ y_{l}=y_{l}(x,\xi),\nu_{l}=\nu_{l}(x,\xi),\\ (x,\xi)\in&\,\mathcal{U}\times(\mathbb{R}^{2}\setminus 0_{2}).\end{split}$
The analog of (5.7) becomes
(9.3) $\begin{split}&\mathcal{G}_{\epsilon}=\text{Op}\big{(}\big{[}\textstyle\sum_{l}\tilde{D}_{l}(x,\xi/|\xi|)+\kappa|\epsilon\xi|^{3}\big{]}^{-1}\big{)}F_{\epsilon},\ x\in\mathcal{U}_{0},\\ &\tilde{D}_{l}(x_{0},\vec{\Theta}_{0}):=2\pi\frac{W^{2}(x_{0},y_{l})|\text{d}_{x}\Phi(x_{0},\alpha_{l})|}{|\Delta_{\Phi}(x_{0},\alpha_{l})|}.\end{split}$
@@
(9.4) $\vec{\Theta}_{0}=\xi_{0}/|\xi_{0}|=\text{d}_{x}H(x_{0})/|\text{d}_{x}H(x_{0})|=\text{d}_{x}\Phi(x_{0},\alpha_{l})/|\text{d}_{x}\Phi(x_{0},\alpha_{l})|$
for all $l$, cf. the convention about the directions of $\text{d}_{x}H$ and $\text{d}_{x}\Phi$ stated above Assumption 3.3. Therefore, we get from (8.14) that the analog of the function $R$ is
@@
Using (4.13) and (6.8), we see that $\Delta\mathcal{G}_{0}(\check{x})$ equals to the sum of the terms analogous to (LABEL:rhs_lim_5):
(9.6) $\begin{split}&\Delta\mathcal{G}_{0}(\check{x})=\Delta f\textstyle\sum_{l}\Upsilon_{l}\big{(}\text{d}_{x}\Phi(x_{0},\alpha_{l})\cdot\check{x}\big{)},\\ &\Upsilon_{l}(r)=2\pi\frac{W^{2}(x_{0},y_{l})}{|\Delta_{\Phi}(x_{0},\alpha_{l})|}\int_{\mathbb{R}}\bar{h}_{l}(u)\\ &\hskip 28.45274pt\times\int_{u}^{u+r}\mathcal{F}^{-1}\left(\big{[}\textstyle\sum_{l}\tilde{D}_{l}(x_{0},\vec{\Theta}_{0})+\kappa|\lambda|^{3}\big{]}^{-1}\right)\bigg{(}\frac{t}{|\text{d}_{x}\Phi(x_{0},\alpha_{l})|}\bigg{)}\text{d}t\text{d}u,\\ &\bar{h}_{l}(u):=\int_{\mathbb{R}}\varphi_{\alpha}(s)\varphi_{p}(\mu s\Phi_{\alpha}^{\prime}(x_{0},\alpha_{l})+u)\text{d}s.\end{split}$